OpenStudy (anonymous):

suppose to solve and round... very lost... do not know where to begin 4^X+1=5^x-1

OpenStudy (anonymous):

\[4^{x+1}=5^{x-1}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

take the log, get \[(x+1)\ln(4)=(x-1)\ln(5)\] then a bunch of algebra to solve for x remember that \(\ln(4)\) and \(\ln(5)\) are just numbers

OpenStudy (anonymous):

got it ?

OpenStudy (anonymous):